OpenStudy (anonymous):

Create a unique example of dividing a polynomial by a monomial and provide the simplified form. Explain, in complete sentences, the two ways used to simplify this expression and how you would check your quotient for accuracy. @ParthKohli @satellite73

Parth (parthkohli):

Erm, why don't you try to create the two of them yourself?

OpenStudy (anonymous):

Because I'm having trouble understanding the lesson as it is. :P

OpenStudy (anonymous):

@phi ? :)

OpenStudy (phi):

a polynomial has multiple terms a monomial (mono means 1 in some language) has 1 term they want you to make up a polynomial, and make up a monomial then divide the monomial into the polynomial

OpenStudy (phi):

see http://www.purplemath.com/modules/polydefs.htm for examples (though you must have seen many examples by now)

OpenStudy (anonymous):

4x^2+6x+2 can be my polynomial, right?

OpenStudy (phi):

sure now make up a monomial I would make it simple like 2x

OpenStudy (anonymous):

Ok, 2x it is,

OpenStudy (anonymous):

@phi http://www.sketchtoy.com/35481798 Did I do it right? :)

OpenStudy (anonymous):

@amistre64 Am I doing this right?

OpenStudy (phi):

almost. the first part is correct but for \[ \frac{6x}{2x} \] the 6/2 is 3 (you got the part right) but x/x is 1 (anything divided by itself is 1) the last term \[ \frac{2}{2x} \] you have to keep the x in the bottom . the 2/2 is 1 so you get \[ \frac{1}{1x} \text{ or just } \frac{1}{x} \]

OpenStudy (phi):

you can check your answer by multiplying it by 2x you should get back the original polynomial

OpenStudy (anonymous):

http://www.sketchtoy.com/35482961 Like this?

OpenStudy (phi):

perfect

OpenStudy (anonymous):

Thanks!

OpenStudy (phi):

no wait!, the 2nd term should not have an x, because x/x is 1

OpenStudy (anonymous):

What should it be?

OpenStudy (phi):

6x/2x = 3

OpenStudy (anonymous):

Okay, I thought so. I fixed it. Thanks again ~